# Active Sparse Bayesian Committee Machine Potential for Isothermal-Isobaric Molecular Dynamics Simulations

### v2

@@ I Introduction
In the last two decades, the application of machine learning potentials (MLPs) in the field of materials simulations has proven to be highly successful, particularly for the accurate prediction of potential energy surfaces (PESs) of the reference ab initio calculations,Unke et al. (2021); Deringer et al. (2021); Bartók et al. (2010); Chen et al. (2023); Gilmer et al. (2017); Batzner et al. (2022); Hajibabaei et al. (2021); Hajibabaei and Kim (2021); Hajibabaei et al. (2021); Myung et al. (2022); Vandermause et al. (2022, 2020); Metcalf et al. (2020); Na and Kim (2021); Zhung et al. (2024); Imbalzano et al. (2021) such as density functional theory (DFT),Laasonen et al. (1993); Richards et al. (2016); Boero et al. (1998) many-body perturbation (MP),Willow et al. (2015); Del Ben et al. (2013, 2015) and coupled-cluster (CC) theories. Moreover, MLPs have an affordable computational cost, allowing one to investigate previously inaccessible chemical phenomena at the first-principles level with currently accessible supercomputers. This breakthrough has paved the way for large-scale and long-time simulations in various scientific domains, including physics, chemistry, biology, and materials science. Several distinct MLP methods have been developed, such as the Behler-Parrinello MLP,Behler and Parrinello (2007); Behler (2011); Pun et al. (2019); Eckhoff and Behler (2021); Schran et al. (2020, 2021); Kapil et al. (2022) Gaussian approximation potentials (GAPs),Bartók et al. (2010, 2013); Bartók and Csányi (2015); Klawohn et al. (2023) sparse Gaussian process regression (SGPR).Hajibabaei et al. (2021); Hajibabaei and Kim (2021); Quiñonero-Candela and Rasmussen (2005); Hajibabaei et al. (2021); Myung et al. (2022), graph neural network potentials (GNNs),Gilmer et al. (2017); Batzner et al. (2022) gradient-domain machine learning (GDML),Chmiela et al. (2017) moment tensor potentials (MTP),Shapeev (2016); Podryabinkin and Shapeev (2017); Novikov et al. (2021) and equivariant message-passing neural networks.Batatia et al. (2022); Batzner et al. (2022); Batatia et al. (2022) The versatility of these methodologies has led to their widespread adoption and successful application across different scientific disciplines, promising valuable insights and outcomes.Lim and Jung (2021)
@@ II Methods
All MD simulations were carried out using the ASE package.Larsen et al. (2017) We employed a time step of 0.5 fs for the integration of the equations of motion that were propagated according to combined Nosé-Hoover and Parrinello-Rahman dynamics.Melchionna et al. (1993); Melchionna (2000); Holian et al. (1990) To acquire the training and testing datasets for analysis, three different systems were chosen. For liquid boron nitride (BN) and LGPS, we trained the BCM potential to the density functional theory (DFT) at the Perdew-Burke-Ernzerhof (PBE) level.Perdew et al. (1996) All the DFT calculations with projector augmented wave (PAW) pseudopotentialsKresse and Joubert (1999) were performed using the Vienna Ab initio Simulation Package (VASP).Kresse and Furthmüller (1996, 1996) We performed isothermal ($NVT$) MD simulations of liquid boron nitride (BN) at temperature of $T=$ 7500 K. The liquid BN system consists of 64 atoms within the cubic box of 7.252 Å. We performed isothermal-isobaric ($NpT$) MD simulations of LGPS at the ambient pressure and temperature ranging from $T=$ 600 to 1000 K. A $\Gamma$-point sampling was employed with an energy cutoff set at 500 eV. We sampled various structures of ice phases (II, III, and V) and liquid water through $NpT$ MD simulations using the MB-pol force fieldBabin et al. (2013, 2014); Riera et al. (2023) at pressures of $p=$ 0.4 and 0.6 GPa, with a temperature of $T=$ 255 K.
@@ III.1 Virial Kernel in Sparse Gaussian Process Regression potential
and $\alpha$ is a hyperparameter.
@@
$=$ $\sum_{i}^{N}\left[\sum_{j}^{m}\dot{\mathcal{K}}_{i}^{\mu}(R,\chi_{j})~{}w_{j}\right]r_{i}^{\nu}.$
The weight vector $\bm{w}$ is obtained by the fitting of $E(R)$, $f_{k}^{\mu}$, and $\mathcal{W}^{\mu\nu}$ to the target energies, forces, virial pressures $\bm{y}_{n}$ for a given dataset $X=\{R_{1},\cdots,R_{n}\}$. The subsequent fitting process of SGPR is nicely reformulated as a form of linear equation with the regularization hyperparameter $\sigma$, Hajibabaei et al. (2021)
$\begin{bmatrix}\bm{K}_{nm}\\ \sigma\bm{L}^{T}\end{bmatrix}\bm{w}_{m}=\begin{bmatrix}\bm{y}_{n}\\ \bm{0}\end{bmatrix},$ (10)
where $\bm{L}$ is the Cholesky factor of $\bm{K}_{mm}$ ($\bm{K}_{mm}=\bm{L}\bm{L}^{T}$, where $\bm{L}$ is lower triangular) and $\bm{0}$ is a columnar vector of zeros with length $m$.
@@ IV.1 Accurate pressure predictions of virial kernel method
In Section III.1, the weight vector $\bm{w}_{m}$ relies on the covariance matrix $\bm{K}_{nm}$ and corresponding target values for energy, forces, virial pressure denoted as $\bm{y}_{n}$. The weight vector’s values vary based on whether the covariance matrix and corresponding target values include the virial pressure term. In this subsection, we compare instantaneous pressure predictions using the weights obtained under two different approaches. This comparative analysis provides insights into the effect of virial kernel in covariance matrix and into the accuracy of pressure prediction. In our subsequent discussion, we will elucidate how the omission of the virial kernel, $\mathcal{W}^{\mu\nu}$, can precipitate abrupt transitions in energy or pressure during the update in active learning. Such discontinuities are problematic in $NpT$ or $NpH$ simulations, undermining its reliability. This phenomenon could raise significant concerns regarding the credibility of simulation results in the absence of virial kernel in SGPR potential.
The first scheme (Scheme 1) involves building $\bm{w}_{m}$ solely with $K_{nm}$ and $\bm{y}_{n}$ that includes energies and forces. The second scheme (Scheme 2) trains $\bm{w}_{m}$ by incorporating virial kernel into $\bm{K}_{nm}$ and $\bm{y}_{n}$ includes energies, forces, and pressures. To substantiate the predictive power of the virial kernel, we evaluated the accuracy of Scheme 1 and Scheme 2 across different phases of materials: high pressure liquid phase BN at $T=$ 7500 K ($NVT$) (Figure S2), solid Li electrolyte LGPS at temperatures ranging from $T=$ 600 to 1000 K ($NpT$), and coexisting phase of ice III and liquid water at pressures of $p=$ 0.4 and 0.6 GPa, with a temperature of $T=$ 255 K ($NpH$). Through these systematic investigations across different systems and phases, we aim to offer a comprehensive evaluation of the virial kernel’s predictive accuracy.
@@
Once we trained the weight vector $\bm{w}_{m}$, we estimated the pressures with the test dataset using Eqs. (8) and (9). This highlights that Scheme 2 is capable of accurately reproducing target instantaneous pressures as opposed to Scheme 1 (Figure 2). BN is an industrially important material which is manufactured at high temperature and pressure. Recently, it was proposed that amorphous BN exhibits low dielectric characteristics and can be used for high performance electronic applications.Hong et al. (2020) Under extreme thermodynamic conditions (T = 7500 K), Scheme 1 struggles to predict the pressure of liquid BN system, whose error is substantially large with the RMSE of 38.43 GPa. However, the inclusion of virial kernel improves its prediction from the RMSE of 38.43 GPa to 6.54 GPa (Figure 2a). The LGPS is one of solid state Li ion electrolytes, and at room temperature, Li ions diffuse while remaining backbone maintains its crystallinity (Figure S3-5). Although the Li ions diffuse away, as it maintains the solid phase, the predictions of both Scheme 1 and Scheme 2 closely align with the reference values with the average predicted pressure value of approximately $p=$ 0 GPa (Figure 2b). Although there was a slight improvement in pressure prediction from the RMSE of 0.26 GPa to 0.13 GPa, the integration of the virial term into the covariance matrix did not manifest a significant effect when the system is in its solid phase under ambient pressure. Under the mild conditions (0.2 GPa $\leq$ $p$leq 0.6 GPa) for the coexisting ice-liquid water system, we again see the pressure prediction improvement by incorporating the virial kernel from RMSE of 1.30 GPa to 0.048 GPa (Figure 2c). Based on the above observations, we conclude that it is essential to incorporate the virial kernel function into covariance matrix, particularly when dealing with a system under high pressure.
IV.2 On-the-flying active learning $NpT$ MD simulations of Liquid Water
@@
On-the-flying active learning $NpT$ MD simulations were performed on liquid water under a pressure of $p=0.2$ GPa and a temperature of $T=270$ K, subsequent to training an expert model through active learning $NVT$ MD simulations at the same temperature. The initial liquid system consisted of 256 water molecules within a cubic box measuring 19.715 Å in box length. Due to its inability to precisely estimate pressure, the $NpT$ MD simulation without the inclusion of virial kernel (Scheme 1) failed to adjust the box’s fluctuations. Updating the dataset $X$, inducing set $z$, and weight vector $\bm{w}$ led to a noticeable sudden change in pressure, disrupting the isenthalipic and isobaric conditions of MD simulations (see Figure 3). On the other hands, when the virial kernel in included in the covariance matrix (Scheme 2), both energy and pressure predictions are smooth during the update in on-the-flying active learning $NpT$ MD simulation, resulting in precise adjustment of the box. In this study, hence, we trained SGPR local expert MLP models for ice II, ice III, and ice V through active learning MD simulations under constant pressure and constant temperature conditions, using an active sparse BCM potential that includes the virial kernel (Scheme 2).
@@
IV.3 Coexistence line of ice-liquid water
@@
We estimate the melting temperatures of ice II, III, and V under pressure conditions of 0.2 $\sim$ 0.6 GPa using the MB-pol water modelRiera et al. (2023). To achieve this, on-the-fly active learning SGPR $NpT$ MD simulations were performed to create BCM MLP for each coexisting systems: ice II-water, ice III-water, and ice V-water. Three separatedly trained SGPR local expert MLP models were then combined into one sparse robust BCM potential. This combined potential was used for $NpH$ MD simulations of coexisting ice-liquid systems. Non-active $NpH$ MD simulations were performed to determine the melting temperatures of ice II, III, and V. Throughout these MD simulations, we maintained the conservation of enthalpy ($H$), while observing fluctuations in the total energy, which is the sum of the potential energy and kinetic energy (see Figure S8).
Figure 4 illustrates the melting points of ice II, III, and V as determined from our $NpH$ MD simulations using the MB-pol water model. The ice-water phase diagram predicted with the MB-pol potential closely align with those obtained from enhanced-coexistence simulations, as reported by Bore and PaesaniBore and Paesani (2023) (see Figure S10). The BCM MLP stabilizes ice V over ice II, consistent with Zhang’s predicted phase diagram.Zhang et al. (2021) The Bore and PaesaniBore and Paesani (2023) using a sophisticated deep neural network potential, however, observed that ice V became a metastable phase, absent from the phase diagram. We attribute this discrepancy to the failure in incorporating high model uncertainty configurations of the water-ice system in the training dataset. In contrast, our active BCM potential adeptly learns model uncertainties on-the-fly, enhancing model accuracy and thereby stabilize ice V in the phase diagram (Figure 4). However, the melting points predicted by the BCM MLP exhibit a slight upward shift in temperature relative to the benchmark MB-pol potential (Figure 4). This deviation primarily stems from the noise present in the energy and force predictions of the BCM potential. Notwithstanding the shifts in melting temperatures, we emphasize the BCM potential’s capability to accurately replicate the physical and thermodynamic properties of the reference MB-pol water model. While the discrepancies in melting temperatures suggest a room for further improvements in accuracy of model, the significance lies in the BCM potential’s reproduction of reference phase diagram. This affirms the BCM potential’s utility in capturing the essence of water’s behavior under various conditions while reducing the computational scaling of training SGPR MLP for large data.

### v3

@@ I Introduction
In the last two decades, the application of machine learning potentials (MLPs) in the field of materials simulations has proven to be highly successful, particularly for the accurate prediction of potential energy surfaces (PESs) of the reference ab initio calculations,Unke et al. (2021); Deringer et al. (2021); Bartók et al. (2010); Chen et al. (2023); Gilmer et al. (2017); Batzner et al. (2022); Hajibabaei et al. (2021); Hajibabaei and Kim (2021); Hajibabaei et al. (2021); Myung et al. (2022); Vandermause et al. (2022, 2020); Metcalf et al. (2020); Na and Kim (2021); Zhung et al. (2024); Imbalzano et al. (2021); Bayerl et al. (2022) such as density functional theory (DFT),Laasonen et al. (1993); Richards et al. (2016); Boero et al. (1998) many-body perturbation (MP),Willow et al. (2015); Del Ben et al. (2013, 2015) and coupled-cluster (CC) theories. Moreover, MLPs have an affordable computational cost, allowing one to investigate previously inaccessible chemical phenomena at the first-principles level with currently accessible supercomputers. This breakthrough has paved the way for large-scale and long-time simulations in various scientific domains, including physics, chemistry, biology, and materials science. Several distinct MLP methods have been developed, such as the Behler-Parrinello MLP,Behler and Parrinello (2007); Behler (2011); Pun et al. (2019); Eckhoff and Behler (2021); Schran et al. (2020, 2021); Kapil et al. (2022) Gaussian approximation potentials (GAPs),Bartók et al. (2010, 2013); Bartók and Csányi (2015); Klawohn et al. (2023) sparse Gaussian process regression (SGPR).Hajibabaei et al. (2021); Hajibabaei and Kim (2021); Quiñonero-Candela and Rasmussen (2005); Hajibabaei et al. (2021); Myung et al. (2022), graph neural network potentials (GNNs),Gilmer et al. (2017); Batzner et al. (2022) gradient-domain machine learning (GDML),Chmiela et al. (2017) moment tensor potentials (MTP),Shapeev (2016); Podryabinkin and Shapeev (2017); Novikov et al. (2021) and equivariant message-passing neural networks.Batzner et al. (2022); Batatia et al. (2022) The versatility of these methodologies has led to their widespread adoption and successful application across different scientific disciplines, promising valuable insights and outcomes.Lim and Jung (2021)
@@ II Methods
All MD simulations were carried out using the ASE package.Larsen et al. (2017) We employed a time step of 0.5 fs for the integration of the equations of motion that were propagated according to combined Nosé-Hoover and Parrinello-Rahman dynamics.Melchionna et al. (1993); Melchionna (2000); Holian et al. (1990) To acquire the training and testing datasets for analysis, three different systems were chosen. For liquid boron Nitride (BN) we trained the BCM potential to DFT at the r2SCAN-D4Furness et al. (2020); Caldeweyher et al. (2017), while for LGPS the potential was trained at the Perdew-Burke-Ernzerhof (PBE) levelPerdew et al. (1996). All the DFT calculations with projector augmented wave (PAW) pseudopotentialsKresse and Joubert (1999) were performed using the Vienna Ab initio Simulation Package (VASP).Kresse and Furthmüller (1996, 1996) We performed isothermal ($NVT$) MD simulations of liquid boron nitride (BN) at temperature of $T=$ 7500 K. The liquid BN system consists of 64 atoms within the cubic box of 7.252 Å. We performed isothermal-isobaric ($NpT$) MD simulations of LGPS at the ambient pressure and temperature ranging from $T=$ 600 to 1200 K. A $\Gamma$-point sampling was employed with an energy cutoff set at 500 eV. We sampled various structures of ice phases (II, III, and V) and liquid water through $NpT$ MD simulations using the MB-pol force fieldBabin et al. (2013, 2014); Riera et al. (2023) at pressures of $p=$ 0.4 and 0.6 GPa, with a temperature of $T=$ 255 K.
@@ III.1 Virial Kernel in Sparse Gaussian Process Regression potential
and $\alpha$ is a hyperparameter, with $\alpha=1$ in our calculations.
@@
$=$ $\sum_{j}^{m}\left[\sum_{i}^{N}\dot{\mathcal{K}}_{i}^{\mu}(R,\chi_{j})r_{i}^{\nu}\right]~{}w_{j}.$
The weight vector $\bm{w}$ is obtained by the fitting of $E(R)$, $f_{i}^{\mu}$, and $\mathcal{W}^{\mu\nu}$ to the target energies, forces, virial pressures $\bm{y}_{n}$ for a given dataset $X=\{R_{1},\cdots,R_{n}\}$. The subsequent fitting process of SGPR is nicely reformulated as a form of linear equation with the regularization hyperparameter $\varepsilon$, Hajibabaei et al. (2021)
$\begin{bmatrix}\bm{K}_{nm}\\ \varepsilon\bm{L}^{T}\end{bmatrix}\bm{w}_{m}=\begin{bmatrix}\bm{y}_{n}\\ \bm{0}\end{bmatrix},$ (10)
where $\bm{L}$ is the Cholesky factor of $\bm{K}_{mm}$ ($\bm{K}_{mm}=\bm{L}\bm{L}^{T}$, where $\bm{L}$ is lower triangular) and $\bm{0}$ is a columnar vector of zeros with length $m$. $\varepsilon$ and other hyperparameters, which are defined in Eq. (10), are optimized by maximizing the log likelihood of $\bm{y}_{n}$.
@@ IV.1 Accurate pressure predictions of virial kernel method
While neural network MLPs can accurately predict virial stress based solely on energies and forces, incorporating virial stress into the training set through a virial stress loss function has been shown to enhance performance by reducing the number of configurations required for training.Zhang et al. (2018) In this work, we demonstrate that including virial stress in the training process is critical for accurate virial stress predictions in SGPR and BCM MLPs. In the SGPR MLP, the training weight parameter vector $\bm{w}_{m}$ relies on the covariance matrix $\bm{K}_{nm}$ and the corresponding target values $\bm{y}_{n}$ for energy, forces, virial pressure in Eq. (10). The values of weight vector, as determined by the solution to Eq. (10), vary depending on whether the virial pressure term is included in the covariance matrix and the corresponding target values. In this subsection, we compare instantaneous pressure predictions using the weights obtained under two different approaches. This comparative analysis provides insights into the effect of virial kernel in covariance matrix and into the accuracy of pressure prediction. In our subsequent discussion, we will elucidate how the omission of the virial kernel, $\mathcal{W}^{\mu\nu}$, can precipitate abrupt transitions in energy or pressure during the update in active learning. Such discontinuities are problematic in $NpT$ or $NpH$ simulations, undermining its reliability. This phenomenon could raise significant concerns regarding the credibility of simulation results in the absence of virial kernel in SGPR potential.
The first scheme (Scheme 1) involves building $\bm{w}_{m}$ solely with $K_{nm}$ and $\bm{y}_{n}$ that includes energies and forces. The second scheme (Scheme 2) trains $\bm{w}_{m}$ by incorporating virial kernel into $\bm{K}_{nm}$ and $\bm{y}_{n}$ includes energies, forces, and pressures. To substantiate the predictive power of the virial kernel, we evaluated the accuracy of Scheme 1 and Scheme 2 across different phases of materials: high-density liquid phase BN at $T=$ 7500 K ($NVT$) (Figure S2), solid Li electrolyte LGPS at temperatures ranging from $T=$ 600 to 1000 K and at atmospheric pressure ($NpT$), and coexisting phase of ice III and liquid water at pressures of $p=$ 0.4 and 0.6 GPa, with a temperature of $T=$ 255 K ($NpH$). Through these systematic investigations across different systems and phases, we aim to offer a comprehensive evaluation of the virial kernel’s predictive accuracy.
@@
Once we trained the weight parameter vector $\bm{w}_{m}$, we estimated the pressures with the test dataset using Eqs. (8) and (9). This highlights that Scheme 2 is capable of accurately reproducing target instantaneous pressures, in contrast to Scheme 1 (Figure 2). BN is an industrially important material which is manufactured at high temperature and pressure. Recently, it was proposed that amorphous BN exhibits low dielectric characteristics and can be used for high performance electronic applications.Hong et al. (2020) Under extreme thermodynamic conditions ($T=$ 7500 K), Scheme 1 struggles to predict the pressure of liquid BN system, resulting in a substantially large error with a root mean square error (RMSE) of 26.4 GPa. However, the inclusion of virial kernel improves the prediction accuracy, reducing the RMSE from 26.4 GPa to 2.8 GPa (Figure 2a).
LGPS is one of the promising candidates for solid state Li-ion electrolytes. At room temperature, LGPS maintains its crystallinity while Li ions diffuse (Figure S4). We find that the pressure predictions for LGPS from both Scheme 1 and Scheme 2 align closely with the reference values because an average predicted pressure is approximately $p=$ 0 GPa (Figure 2b). The integration of the virial term into the covariance matrix does not significantly affect the pressure when the system is in its solid phase at ambient pressure. Still, there is a slight improvement in prediction accuracy, with an RMSE decreasing from 0.26 GPa to 0.13 GPa.
Under the mild conditions (0.2 GPa $\leq$ $p$leq 0.6 GPa) of the coexisting ice-liquid water system, incorporating the virial kernel significantly improved the pressure prediction from an RMSE of 1.30 GPa to 0.048 GPa (Figure 2c). These results demonstrate that including the virial kernel function in the covariance matrix is crucial for SGPR MLP, especially when dealing with high-pressure systems.
IV.2 On-the-flying active learning $NpT$ MD simulations of LGPS
In 2012, Mo et. al. reported that $NVT$ first-principles MD simulations of LGPS, conducted from 600 K to 1500 K, indicated no breaking of P-S or Ge-S bonds.Mo et al. (2012) They concluded that LGPS is a promising and structurally stable material for all-solid-state batteries. However, at the extremely high temperature of 1500 K, the system’s pressure in a fixed volume is expected to be exceedingly high. To address this issue, on-the-flying active learning $NpT$ MD simulations were performed under atmospheric pressure ($\sim$0.101 MPa) at various temperatures for solid Li electrolyte LGPS, with each run lasting over 100 ps (Figure 3 (a)).
The average potential energy of LGPS (Figure 3 (b)) shows two distinct linear regions: one for the solid phase (600$\sim$900 K) and another for the liquid phase (1100$\sim$1200 K). These observations from both the mean square displacement (MSD) of Li in LGPS and potential energy analyses provide clear evidence for a phase transition in LGPS around 1000 K, transitioning from a solid to a liquid state. The MSD demonstrates a clear temperature-dependent behavior, with higher temperatures resulting in increased Li ion mobility. At lower temperatures (600$\sim$900 K), the MSD shows a gradual increase, indicating solid-state diffusion. However, above 1000 K, the slope and magnitude of the MSD increase rapidly, especially at 1100 K and 1200 K. The rapid change in Li ion mobility is consistent with the potential energy discontinuity observed in Figure 3 (b), supporting a phase transition from a solid to a liquid state. This transition significantly enhances Li ion mobility, as evidenced by the steep increase in MSD values at higher temperatures.
This observation aligns well with the experimental data,Hori et al. (2015) wherein differential thermal analysis (DTA) was used to elucidate the phase diagram for the quasi-binary system between Li4GeS4 and Li3PS4. The experimental investigation revealed that LGPS exhibits an incongruent melting point at 650 C (923 K). We confirmed that the activation energy $E_{a}$ for Li diffusion in LGPS in its solid state (below 1000 K) is about 0.18 eV. This value is similar to the previously reported value of 0.17 eV.Hajibabaei and Kim (2021)
Previous computational studies on various materials, like MgO, have shown similar trends in potential energy changes with temperature. These studies using the deep neural network based MLPs have found a sudden shifts in energy due to phase transition.Wisesa et al. (2023, 2023); Andolina and Saidi (2023)
Figure 3: (a) Mean square displacement (MSD) of Li in LGPS at various temperatures (600 K: blue, 700 K: orange, 800 K: green, 900 K: red, 1000 K: purple, 1100 K: brown, 1200 K: pink). (b) Average potential energy of LGPS during $NpT$ simulations. The system undergoes phase transition from a solid to a liquid state above 1000 K.
IV.3 On-the-flying active learning $NpT$ MD simulations of Liquid Water
@@
On-the-flying active learning $NpT$ MD simulations were performed on liquid water under a pressure of $p=0.2$ GPa and a temperature of $T=270$ K, subsequent to training an expert model through active learning $NVT$ MD simulations at the same temperature. The initial liquid system consisted of 256 water molecules within a cubic box measuring 19.715 Å in box length. Due to its inability to precisely estimate pressure, the $NpT$ MD simulation without the inclusion of virial kernel (Scheme 1) failed to adjust the box’s fluctuations. Updating the dataset $X$, inducing set $z$, and weight vector $\bm{w}$ led to a noticeable sudden change in pressure, disrupting the isenthalipic and isobaric conditions of MD simulations (see Figure 4). On the other hands, when the virial kernel is included in the covariance matrix (Scheme 2), both energy and pressure predictions are smooth during the update in on-the-flying active learning $NpT$ MD simulation, resulting in precise adjustment of the box. In this study, hence, we trained SGPR local expert MLP models for ice II, ice III, and ice V through active learning MD simulations under constant pressure and constant temperature conditions, using an active sparse BCM potential that includes the virial kernel (Scheme 2).
@@
IV.4 Coexistence line of ice-liquid water
@@
We estimate the melting temperatures of ice II, III, and V under pressure conditions of 0.2 $\sim$ 0.6 GPa using the MB-pol water model.Riera et al. (2023) To achieve this, we performed on-the-fly active learning SGPR $NpT$ MD simulations to create a BCM MLP for each coexisting system: ice II/water, ice III/water, and ice V/water. Three separately trained SGPR local expert MLP models were then combined into one sparse robust BCM potential. This combined potential was used for non-active learning $NpH$ MD simulations of coexisting ice-liquid systems. No updates were made to the inducing sets and datasets in the non-active MD simulations. Non-active $NpH$ MD simulations were performed to determine the melting temperatures of ice II, III, and V. Throughout these MD simulations, we maintained the conservation of enthalpy ($H$), while observing fluctuations in the total energy, which is the sum of the potential energy and kinetic energy (Figure S6).
Figure 5 illustrates the melting points of ice II, III, and V as determined from our $NpH$ MD simulations using the MB-pol water model. The ice-water phase diagram predicted with the MB-pol potential closely align with those obtained from enhanced-coexistence simulations, as reported by Bore and PaesaniBore and Paesani (2023) (see Figure S8). The BCM MLP stabilizes ice V over ice II, consistent with Zhang’s predicted phase diagram.Zhang et al. (2021) The Bore and PaesaniBore and Paesani (2023) using a sophisticated deep neural network potential, however, observed that ice V became a metastable phase, absent from the phase diagram. We attribute this discrepancy to the failure in incorporating high model uncertainty configurations of the water-ice system in the training dataset. In contrast, our active BCM potential adeptly learns model uncertainties on-the-fly, enhancing model accuracy and thereby stabilize ice V in the phase diagram (Figure 5). However, the melting points predicted by the BCM MLP exhibit a slight upward shift in temperature relative to the benchmark MB-pol potential (Figure 5). This deviation primarily stems from the noise present in the energy and force predictions of the BCM potential. Notwithstanding the shifts in melting temperatures, we emphasize the BCM potential’s capability to accurately replicate the physical and thermodynamic properties of the reference MB-pol water model. While the discrepancies in melting temperatures suggest a room for further improvements in accuracy of model, the significance lies in the BCM potential’s reproduction of the reference phase diagram. This affirms the BCM potential’s utility in capturing the essence of water’s behavior under various conditions while reducing the computational effort required to train SGPR MLP on large datasets.